OpenStudy (anonymous):

Solve the following equation for z: 15 = 10z/[2-(7x4)]

OpenStudy (australopithecus):

Do you know what to do first?

OpenStudy (australopithecus):

You are simply isolating the exponent z on one side of the equation

OpenStudy (anonymous):

no, this is from my general chem lab 1 quiz

OpenStudy (australopithecus):

Really? Because this is just basic algebra, but you do need to have a strong grasp of algebra to be a chemist so meh smeh First step, is getting rid of what is in the denominator on the right, do you know how you could go about that? 15 =10z/[2-(7x4)]

OpenStudy (australopithecus):

by getting rid of it I mean bringing it to the left side of the equation and just leaving 10z on the right side

OpenStudy (anonymous):

I haven't did that in some years. I appreciate the help though

OpenStudy (australopithecus):

Equal signs simply means that the expressions on both sides of it are the same thing (they are equal). As long as you do to one side what you do to the other the equality is maintained, which means you can divide, multiply, subtract etc one side of an equality by anything as long as you do it to the other side. For instance 15x = 10y/(3-x) if we wanted to isolate y it would be as simple as first realizing that you can if you divide something by itself it is equal to 1 and anything multiplied by 1 is itself. for example 4x/4x = 1 (2-3x)/(2-3x) = 1 (3333+z+y+z)/(3333+z+y+z) = 1 so we can apply this rule to the above equation, (3-x)*15x = (3-x)*(10y/(3-x)) = (3-x)*15x = (3-x)(10y)/(3-x) (3-x)*15x = 10y*1 (3-x)*15x = 10y

OpenStudy (australopithecus):

do you follow?

OpenStudy (anonymous):

I understand

OpenStudy (australopithecus):

so can you solve your problem now?

OpenStudy (anonymous):

no lol, im still alil lost but i understand alil

OpenStudy (australopithecus):

well can you do your first step?

OpenStudy (australopithecus):

now

OpenStudy (australopithecus):

The step I mentioned above, I dont mind walking you through this

OpenStudy (anonymous):

would it be (2-{7x4)]15=10z

OpenStudy (australopithecus):

YES! Now ask yourself how can you get rid of the 10 from the right side of the equation

OpenStudy (anonymous):

divide both sides by 10

OpenStudy (australopithecus):

Right!

OpenStudy (australopithecus):

Problem solved

OpenStudy (australopithecus):

This stuff is easy right

OpenStudy (anonymous):

whats the answer lol, its just gonna be [2-(7x4)] over 15

OpenStudy (australopithecus):

write out what you think the answer is?

OpenStudy (australopithecus):

you just need to divide both sides by 10 and you are done

OpenStudy (australopithecus):

All you have to do is isolate z on one side of the equation, can you show me your final answer?

OpenStudy (australopithecus):

your answer is wrong btw, You are at this step, (2-{7x4)]15=10z you said divide both sides by 10 makes sense, (2-{7x4)]15/10=10z/10 (2-{7x4)]15/10=(1)z (2-{7x4)]15=z That is your answer

OpenStudy (australopithecus):

make sure you understand what I did this stuff is fundamental and wont ever go away if you plan on doing any technical job

OpenStudy (australopithecus):

although maybe not idk

OpenStudy (australopithecus):

oops, (2-{7x4)]15/10=z is the final answer forgive me

OpenStudy (australopithecus):

I hope you understand